Corey is 58 years younger than Catherine. 9 years ago, Catherine's age was 3 times Corey's age. How old is Corey now
Answer by ikleyn(52787)

.

Let Corey is x years old now.

Then Catherine's age is (x+58) years.


9 years ago . . . 


    (x+58)-9 = 3*(x-9)


Simplify and solve


    x + 49 = 3x - 27

    49 + 27 = 3x - x

    76      = 2x

     x      = 76/2 = 38.


ANSWER.  Corey now is 38 years old.